# Open Stamped Parts Dataset*

### v2

@@ Abstract
We present the Open Stamped Parts Dataset (OSPD), featuring synthetic and real images of stamped metal sheets for auto manufacturing. The real part images, captured from 7 cameras, consist of 7,980 unlabeled images and 1,680 labeled images. In addition, we have compiled a defect dataset by overlaying synthetically generated masks on 10% of the holes. The synthetic dataset replicates the real manufacturing environment in terms of lighting and part placement relative to the cameras. The synthetic data includes 7,980 training images, 1,680 validation images and 1,680 test images, each with bounding box and segmentation mask annotations around all holes. 10% of the holes in the synthetic data mimic defects generated in the real image dataset. We trained a hole-detection model on the synthetic-OSPD, achieving a modified recall score of 67.2% and a precision of 94.4% . We anticipate researchers in the auto manufacturing and broader machine learning and computer vision communities using OSPD to advance the state of the art in defect detection of stamped holes in the metal-sheet stamping process. The dataset is available for download at: https://tinyurl.com/hm6xatd7.
@@ I INTRODUCTION
Manual inspection of stamped parts is time-consuming and susceptible to human errors. During this process, inspectors remove sheets from the production line to individually check defects, for example, missing stamped holes, a laborious task due to the numerous, sometimes very small, holes. Consequently, manual inspection of every sheet is impractical. The high cost of manual inspection coupled with the costly consequence of defects, have spurred research interest in automated defect detection systems. Automation promises increased reliability, scalability, and cost-effectiveness.
@@
Addressing the data challenge could involve training object detection models using synthetically generated images. Still, replicating the manufacturing environment proves difficult, often leading to a domain gap between synthetic and real data [8]. In our dataset, one such gap lies in the background of the stamped sheet: synthetic images feature empty backgrounds, while real images depict conveyors and wire meshes. Refer to Fig. 1 for a visual representation of this domain gap in our data. Recent strides in generative deep learning techniques offer promise in minimizing this gap, enhancing the robustness of object detection solutions [9, 10].
Figure 1: a. Example of synthetically rendered stamped metal sheet b. Example of real stamped sheet metal on conveyor belt in factory setting.
@@
Our hope is that the proposed baseline model, evaluation metrics, and datasets (both real and synthetic) of sheet metal images, the Open Stamped Parts Dataset (OSPD), will facilitate further development of efficient and reliable machine vision models to automate the process of defect inspection in auto manufacturing.
@@
  • A labeled set of grayscale real images with bounding box annotations  [^1] Metadata includes camera name and trigger ${}^{,}$ [^2] See Fig. 2 for diagram of camera fixture. Trigger distance (measured in millimeters) measures the distance the sheet metal has travelled along the conveyor relative to the fixture ${}^{,}$ [^3] Metadata includes whether or not the hole is graded. A graded hole is defined as one where all edges of the hole are visible in the reference image and the hole size is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (also referred to as a “view”) where there is neither sheet metal part shift relative to the camera nor camera shift relative to its expected position..
  • A labeled set of grayscale real images with simulated missing hole defects and bounding box annotations  [^1] Metadata includes camera name and trigger ${}^{,}$ [^2] See Fig. 2 for diagram of camera fixture. Trigger distance (measured in millimeters) measures the distance the sheet metal has travelled along the conveyor relative to the fixture ${}^{,}$ [^3] Metadata includes whether or not the hole is graded. A graded hole is defined as one where all edges of the hole are visible in the reference image and the hole size is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (also referred to as a “view”) where there is neither sheet metal part shift relative to the camera nor camera shift relative to its expected ${}^{,}$ [^4] Metadata includes whether or not the hole is masked..
  • A large scale, labeled synthetic dataset to mimic the real data with segmentation masks and bounding boxes [^1] Metadata includes camera name and trigger ${}^{,}$ [^3] Metadata includes whether or not the hole is graded. A graded hole is defined as one where all edges of the hole are visible in the reference image and the hole size is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (also referred to as a “view”) where there is neither sheet metal part shift relative to the camera nor camera shift relative to its expected ${}^{,}$ [^5] Metadata includes the angle between the relative normal of the hole center and the camera..
@@ II-B Surface Inspection Datasets
Dataset Image Type #Images Task Synth. / Real
PCB [12]
HDR RGB
1,386
Classification Detection Registration
Synth.
Severstal [13]
grayscale
18,074
Semantic Segmentation
Real
DAGM [14]
grayscale
16,100
Detection
Synth.
NEU [15]
grayscale
1,800
Detection
Real
OSPD
grayscale
11,340
Detection Instance Segmentation
Synth.
OSPD
grayscale
9,660
Detection
Real
@@ II-C Transfer Learning Between Synthetic and Real Data
When a new part is onboarded, little to no real data is available for training a neural network. Training on synthetically generated data becomes imperative in these scenarios, and several works aim to bridge the gap between the real and synthetic data through different techniques [9] [10]. See the Experiments section for a performance comparison of models trained on synthetic data that has and has not been transformed by an image-to-image translation model.
@@ III-A Camera Fixture Setup
The OSPD’s image collection involves a camera fixture equipped with 10 cameras positioned across its left, right, entrance, and exit sides, totaling 40 cameras. Entrance/exit designations are relative to conveyor belt movement, while right/left designations are relative to viewing from the entrance side. Each camera is named “ZxCAMy,” where x=1 and x=2 specify left and right-side camera respectively and $y\in[1,20]$. Refer to Fig. 2 for a diagram illustrating the fixture, annotated with camera names. In the OSPD, images are exclusively sourced from a subset of left-sided cameras. Each camera may trigger multiple times at varying distances as the part moves through the fixture. In OSPD, the COCO-formatted metadata for each image includes the camera name and trigger distance within the “cam_name” and “trigger_mm” fields, respectively. Lights on the fixture are infrared and cameras have infrared filters installed on their lens’ to block out other light sources.
Figure 2: a. Subset of cameras labeled on fixture’s entrance side. Exit side mirrors entrance side, but with even-numbered cameras b. Illustration of conveyor passing through fixture.
@@ III-B Data Generation and Acquisition
The synthetic data was acquired using Unreal Engine [16], by carefully configuring parameters such as camera angle, focal length, and lighting to replicate the actual production environment. The simulation generated images of a stamped sheet with 7 distinct hole categories. To mimic missing stamped-hole defects, we applied a mask to a hole in any image with 10% probability. It’s important to note that the conveyor and other background elements present in the manufacturing facility were not generated in Unreal Engine, resulting in a domain gap between synthetic and real data, which we aim to address.
The real data used for testing was taken over the course of two days in March and April, 2023. To assess a model’s ability to identify missing holes accurately, while lacking actual instances of missing holes in real data, we transformed the test dataset by applying masks to a subset of the holes using an inpainting technique [17]. Refer to figure 3 for representations of two different masked holes.
Figure 3: Examples of masked holes. a. Hole category, DY1, outlined in red as masked b. Hole category, DY2, outlined in red as masked.
@@
The unlabeled real data was collected over a period of 12 inspections days in August and September, 2023. Inspections were taken by the same cameras at the same trigger distances as the inspections in the labeled real data. Unlabeled real data is provided in the same quantity as the synthetic training set: there are 380 instances for each of the 21 views for a total of 7,980 images.
@@ III-C Labeling Process
Labelers were instructed to create bounding boxes around entire holes, despite potential difficulties in detection, given prior knowledge that all holes were present during inspections. Examples of challenging-to-detect holes, appropriately annotated, are depicted in Fig 4. For testing purposes, only graded holes were considered. See footnote 3 on page 2 for the definition of a graded hole. Additional details regarding evaluation metrics are available in the Metric section.
Figure 4: Examples of labeled hole images. Each hole is enclosed in a bounding box colored according to its category ID. Arrows point to the referenced hole annotations a. Small holes, difficult to detect by the human eye, were labeled b. Partially visible, low contrast holes were labeled.
@@ IV IMAGE SETS & STATISTICS
Our primary objective is to accurately detect the presence of all holes, and thereby identify any missing ones. In practical production scenarios, the occurrence of a missing hole is extremely rare (<1% chance); however, the consequences of not detecting a missing hole are high.
Considering that neural networks learn from examples, and noting that there are limited instances of missing holes in actual production data, we opted to increase the probability of missing holes in the synthetic training data beyond the expected frequency in real-world scenarios.
@@ V-A Object Detection
YOLO, a leading CNN based object detection model serves as our baseline model [4]. Specifically, we train a YOLOv7 model [^1] Code available at https://github.com/WongKinYiu/yolov7, as it demonstrates superior speed and accuracy compared to prior YOLO architectures [19].
The YOLOv7 model, starting with pre-trained model weights and p5 hyperparameters, was trained for 1,000 epochs with a batch size of 8 on an NVIDIA A100 GPU. Training images were resized to 640x640, padding the smaller dimension of the original image to maintain the aspect ratio.
@@
    If $a_{r}<=1,100$, filter out all $i_{p}$ except for $i_{p}$ whose center is nearest to $f_{r}$. [^3] Manhattan distance is used for calculating distance [23]. After all $f_{r}$ have been considered, each $f_{r}$ will be assigned to zero, one, or many $i_{p}$.
@@
    Find the maximum cardinality matching by solving the minimum cost assignment algorithm where cost values in our cost matrix are defined as the Manhattan distance between the center of $i_{p}$ and $f_{r}$ and the boolean matrix is derived from the previous step where we found potential matches [23]. In this way, each $i_{p}$ is assigned to at most one $f_{r}$ and each $f_{r}$ is assigned to at most one $i_{p}$ [24].
@@ V-C Metric
[…]
@@ VI EXPERIMENTS
[…]
TABLE II: Model Results
[…]
Synthetic – multi-class
60
67.2
94.4
79.9
I2I – multi-class
65
58.7
73.1
65.9
Synthetic – binary + custom category ID protocol
75
61.1
81.6
70.9
I2I – binary + custom category ID protocol
65
57.7
86.4
70.7
VII CONCLUSION
We introduced OSPD, a dataset comprising synthetic and real stamped metal sheets, which aims to enhance defect detection of stamped holes in manufacturing. Our custom metric prioritizes identifying missing holes, and we presented a baseline model as a starting point for future enhancements. Finally, our statistical analysis of the data offers contextual insights and identifies potential sources of model bias.
@@ ACKNOWLEDGMENT
Thank you to Anishi Mehta for her contributions in literature review, editing and planning. Thank you to Rosemary Binny, Sherry Courington, Carlos Cardenas, Beatrice Lovely, Anishi Mehta, Joseph Muniz Jr., and Jose Rangel, who labeled our data. Thank you to Hilton Shumway, Sean Sloan and Anthony Swink who worked with Unreal Engine to generate synthetic data. Thank you to Jose Armando Lopez Jr., Steve Rangos, Thomas Sandrisser, and Tyler Wingo for their review and edits. Thank you to Jose Armando Lopez Jr. and Thomas Sandrisser for their support of this project.

### v3

@@ Abstract
We present the Open Stamped Parts Dataset (OSPD), featuring synthetic and real images of stamped metal sheets for auto manufacturing. The real part images, captured from 7 cameras, consist of 7,980 unlabeled images and 1,680 labeled images. In addition, we have compiled a defect dataset by overlaying synthetically generated masks on 10% of the holes. The synthetic dataset replicates the real manufacturing environment in terms of lighting and part placement relative to the cameras. The synthetic data includes 7,980 training images, 1,680 validation images and 1,680 test images, each with bounding box and segmentation mask annotations around all holes. 10% of the holes in the synthetic data mimic defects generated in the real image dataset. We trained a hole-detection model on the synthetic-OSPD, achieving a modified recall score of 67.2% and a precision of 94.4% . We anticipate researchers in auto manufacturing use OSPD to advance the state of the art in defect detection of stamped holes in the metal-sheet stamping process. The dataset is available for download at: https://tinyurl.com/hm6xatd7.
@@ I INTRODUCTION
Manual inspection of stamped parts is time-consuming and susceptible to human errors. During this process, inspectors remove sheets from the production line to individually check for defects such as missing stamped holes. This is a laborious task due to the numerous, sometimes very small, holes. Consequently, manual inspection of every sheet is impractical. The high cost of manual inspection coupled with the costly consequence of defects, have spurred research interest in automated defect detection systems. Automation promises increased reliability, scalability, and cost-effectiveness.
@@
Addressing the data challenge could involve training object detection models using synthetically generated images. Still, replicating the manufacturing environment proves difficult, often leading to a domain gap between synthetic and real data [8]. In our dataset, one such gap lies in the background of the stamped sheet: synthetic images have no background artifacts, while real images depict conveyors and wire meshes. Refer to Fig. 1 for a visual representation of this domain gap in OSPD. Recent strides in generative deep learning techniques offer promise in minimizing this gap, enhancing the robustness of object detection solutions [9, 10].
Figure 1: Example of synthetically rendered (a.) and real stamped sheet metal (b.)
@@
Our hope is that the proposed baseline model, evaluation metrics, and datasets (both real and synthetic) of sheet metal images, the Open Stamped Parts Dataset (OSPD), will facilitate further development of efficient and reliable machine vision models to automate the process of defect inspection in auto manufacturing. Moreover, the provided datasets can serve as a playground for testing ideas which can then be generalized across different factory settings and manufactured parts.
@@
  • A labeled set of grayscale real images with bounding box annotations  [^1] Metadata includes camera name and trigger distance., [^2] See Fig. 2 for camera fixture diagram. Trigger distance (measured in millimeters) is the distance the sheet metal travels along the conveyor relative to the fixture entrance., [^3] Metadata includes whether the hole is graded. A graded hole has all hole edges visible in the reference image and is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (a “view”) without any sheet metal shift or camera shift relative to the expected positions..
  • A labeled set of grayscale real images with simulated missing hole defects and bounding box annotations  [^1] Metadata includes camera name and trigger distance., [^2] See Fig. 2 for camera fixture diagram. Trigger distance (measured in millimeters) is the distance the sheet metal travels along the conveyor relative to the fixture entrance., [^3] Metadata includes whether the hole is graded. A graded hole has all hole edges visible in the reference image and is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (a “view”) without any sheet metal shift or camera shift relative to the expected positions., [^4] Metadata includes whether or not the hole is masked..
  • A large scale, labeled synthetic dataset to mimic the real data with segmentation masks and bounding boxes [^1] Metadata includes camera name and trigger distance., [^3] Metadata includes whether the hole is graded. A graded hole has all hole edges visible in the reference image and is 100 pixels or larger. A reference image is a single real image for each camera and trigger distance combination (a “view”) without any sheet metal shift or camera shift relative to the expected positions., [^5] Metadata includes the angle between the hole center’s relative normal and the camera..
@@ II-B Surface Inspection Datasets
Dataset Image Type #Images Task Synth. / Real Defect Classes
PCB [12] HDR RGB 1,386 Classification, Detection, Registration Synth. open circuit, short, spurious copper, missing hole, mouse bite, spur
Severstal [13] grayscale 18,074 Semantic Segmentation Real pitted, crazing, scratches, patches
DAGM [14] grayscale 16,100 Detection Synth. texture anomaly
NEU [15] grayscale 1,800 Detection Real rolled-in scale, patches, crazing, pitted surface, inclusion, scratches
OSPD grayscale 11,340 Detection, Instance Segmentation Synth. missing hole
OSPD grayscale 9,660 Detection Real missing hole
@@ II-C Transfer Learning Between Synthetic and Real Data
When a new part is onboarded, little to no real data is available for training a neural network. Training on synthetically generated data becomes imperative in these scenarios, and several works aim to bridge the gap between the real and synthetic data through different techniques [9], [10]. See the Experiments section for a performance comparison of models trained on synthetic data that has and has not been transformed by an image-to-image (I2I) translation model.
@@ III-A Camera Fixture Setup
The OSPD’s image collection uses a fixture with 40 cameras positioned across the left, right, entrance, and exit sides. 7 left-sided cameras took 21 unique captures in the OSPD. Entrance/exit designations are relative to conveyor belt movement, while right/left are relative to viewing from the entrance side. Cameras are named “ZxCAMy,” where x=1 and x=2 specify left and right-side camera respectively and $y\in[1,20]$. Refer to Fig. 2 for fixture diagram with camera names. Each camera may trigger multiple times as the part moves through the fixture. OSPD metadata includes camera name (“cam_name”) and trigger distance (“trigger_mm”). Lights on the fixture are infrared and cameras have infrared filters to block other light sources.
Figure 2: a. Entrance side fixture with subset of cameras labeled. Exit side mirrors entrance side, but with even-numbered cameras b. Illustration of conveyor passing through fixture.
@@ III-B Data Generation and Acquisition
The synthetic data was acquired using Unreal Engine [16], by carefully configuring parameters such as camera angle, focal length, and lighting to replicate the actual production environment. Small amounts of random camera shift, part shift, and lighting changes were included to create synthetic data to simulate changes in production conditions. The simulation generated images of a stamped sheet with 7 distinct hole categories. To mimic missing stamped-hole defects, we applied a mask to a hole in any image with 10% probability. We note that the conveyor and other background elements present in the manufacturing facility were not generated in Unreal Engine, resulting in a domain gap between synthetic and real data, which we aim to address.
The real data used for testing was taken over the course of two days in March and April, 2023. To assess a model’s ability to identify missing holes accurately, while lacking actual instances of missing holes in real data, we transformed the test dataset by applying masks to a subset of the holes using an inpainting technique [17]. Refer to Fig. 3 for representations of two different masked holes.
Figure 3: Examples of masked holes, hole category DY1 (a.) and hole category DY2 (b.)
@@
Unlabeled real data was collected over 12 inspections days in August and September, 2023. Inspections were taken by the same cameras and trigger distances as the inspections in the labeled real data. Unlabeled real data is provided in the same quantity as the synthetic training set: 380 instances for each of the 21 views for a total of 7,980 images.
@@ III-C Labeling Process
Labelers were instructed to create bounding boxes around entire holes, despite potential difficulties in detection, given prior knowledge that all holes were present during inspections. Examples of challenging-to-detect holes, appropriately annotated, are depicted in Fig. 4. For testing purposes, only graded holes were considered. See footnote 3 on page 2 for the definition of a graded hole. Additional details regarding evaluation metrics are available in the Metric section.
Figure 4: Examples of labeled hole images, small holes (a.) and partially visible, low contrast holes (b.) Arrows point to the referenced hole annotations.
@@ IV IMAGE SETS & STATISTICS
Since missing holes have no distinctive features, we invert our problem: we formulate the task as detection of holes and thereby, we infer which holes are missing. In practical production scenarios, the occurrence of a missing hole is extremely rare (<1% chance); however, the consequences of not detecting a missing hole are high.
@@ V-A Object Detection
YOLO, a leading CNN based object detection model serves as our baseline model [4]. Specifically, we train a YOLOv7 model [^1] Code available at https://github.com/WongKinYiu/yolov7, as it demonstrates superior speed and accuracy compared to prior YOLO architectures [19]. The YOLOv7 model was trained for 1,000 epochs with a batch size of 8 on an NVIDIA A100 GPU. Training images were resized to 640x640, padding the smaller dimension of the original image to maintain the aspect ratio.
@@
    If $a_{r}<=1,100$, filter out all $i_{p}$ except for $i_{p}$ whose center is nearest to $f_{r}$. [^3] Euclidean distance is used for calculating distance [23]. After all $f_{r}$ have been considered, each $f_{r}$ will be assigned to zero, one, or many $i_{p}$.
@@
    Find the maximum cardinality matching by solving the minimum cost assignment algorithm where cost values in our cost matrix are defined as the Euclidean distance between the center of $i_{p}$ and $f_{r}$ and the boolean matrix is derived from the previous step where we found potential matches. In this way, each $i_{p}$ is assigned to at most one $f_{r}$ and each $f_{r}$ is assigned to at most one $i_{p}$ [24].
@@ V-C Metric
Our foremost goal is to accurately identify missing holes; classifying a missing hole as found can significantly compromise this objective. Misclassification can occur in two ways: the model either recognizes a missing hole as a hole or identifies a different object as a hole. Consequently, our metric aims to reflect our strong preference for zero false positives in both of these scenarios. Provided there is a low occurrence of false positives on objects, our secondary objective involves minimizing instances where the model incorrectly labels an existing hole as missing. As such, we define a best model as one with the highest Fβ score:
@@
TABLE II: Model Results
Model Conf. Thresh. AR AMR AP F0.9 Score Modified F0.9 Score
Synthetic – multi-class 60 67.2 67.2 94.4 79.9 79.9
I2I – multi-class 65 61.6 58.7 73.1 67.5 65.9
Synthetic – binary + custom category ID protocol 75 61.1 61.1 81.6 61.1 70.9
I2I – binary + custom category ID protocol 65 60.3 57.7 86.4 72.4 70.7
@@ VI EXPERIMENTS
Table II displays the confidence threshold, average recall (AR), average modified recall (AMR), and average precision (AP) that yields the highest Modfied F0.9 score where confidence threshold was tested at $5\%$ increments from $50\%$ to $95\%$. Synthetic models are YOLOv7 models trained on data generated directly by Unreal Engine while I2I models are YOLOv7 models trained on data generated by Unreal Engine and then fed through a CUT image-to-image translation model [27]. Multi-class models predict each object as one of several hole classes and are fed through the filtering method outlined in the Model Description section for increasing precision. The binary + custom category ID protocol models predict each object under a single hole class and are fed through the protocol outlined in the Model Description section for filtering and category assignment. Because we are interested in optimizing the modified recall metric, we chose to evaluate each model using the model checkpoint which yields the highest recall on the validation set during training. Training a multi-class model with purely synthetic data yields the highest modified recall, 67.2, and highest precision, 94.4. I2I appears to add no increase in performance. One reason for this may be because we approached I2I in a simplistic way: splitting image into 16x16 patches, passing through I2I model, and stitching patches back together.
VII FUTURE RESEARCH
In future work one could expand the synthetic and real dataset to include more types of stamping defects. One could investigate more advanced I2I techniques or background subtraction models to remove background distractions present in real data, bridging the domain gap between synthetic and real data. This may also improve the generalization capabilities of the defect detection model across different factory settings.
@@ ACKNOWLEDGMENT
Thank you to Jose Armando Lopez Jr. and Thomas Sandrisser for their support of this project. We also acknowledge support from Rosemary Binny, Sherry Courington, Carlos Cardenas, Beatrice Lovely, Anishi Mehta, Joseph Muniz Jr., Jose Rangel, Hilton Shumway, Sean Sloan and Anthony Swink, Tyler Wingo and Steve Rangos.